### v1

# A Bayesian Dirichlet Auto-Regressive Conditional Heteroskedasticity Model for Compositional Time Series

Abstract
We analyze a high-frequency, real-world compositional dataset: Airbnb’s daily currency-fee proportions across multiple regions. These proportions sum to one and exhibit changing levels of volatility, particularly following disruptions such as the COVID-19 pandemic. Standard compositional models often assume constant variance, but our data reveal episodes of elevated uncertainty and clustering in volatility.
Motivated by these empirical patterns, we propose a Bayesian Dirichlet Auto-Regressive Moving Average framework extended with a Dirichlet Auto-Regressive Conditional Heteroskedasticity (DARCH) component—the B-DARMA-DARCH model. This novel approach retains the Dirichlet likelihood, ensuring valid predictions on the simplex, and introduces a time-varying precision parameter via an ARMA recursion. As a result, B-DARMA-DARCH captures both mean dynamics and stochastic volatility in compositional time series.
Simulations and empirical tests on the Airbnb currency data show that B-DARMA-DARCH outperforms simpler Dirichlet ARMA and Bayesian log-ratio VARMA models in forecasting accuracy, residual diagnostics, and interval coverage. By accommodating volatility clustering and structural breaks, B-DARMA-DARCH provides a robust framework for compositional data analysis in finance and beyond, supporting improved risk management and uncertainty quantification in settings where proportions and their volatility are of primary interest.
Keywords: Airbnb, Additive Log Ratio, Bayesian multivariate time series, Compositional time series, Currency Volatility, Data Science, Dirichlet distribution, Finance, GARCH, Generalized ARMA model, Heteroskedasticity, Hospitality Industry, Risk Management, Simplex, Vector ARMA model, Volatility Clustering
1 Introduction
Airbnb operates a two-sided marketplace with millions of listings worldwide, and it accommodates transactions in over 100 currencies. While individual hosts set their own nightly rates, the platform must still forecast these prices to predict total daily revenue. Complicating matters, travelers can pay in a currency of their choosing (often their local currency), resulting in a dynamic compositional time series of daily fee shares across USD, EUR, CAD, and other currencies. This currency composition is crucial for accurately converting booked stays into a consolidated U.S. Dollar (USD) amount—a requirement for company-wide financial reporting, treasury operations, and broader risk management.
To forecast revenue in USD, one must know (i) the nightly prices that hosts set in their local currency, and (ii) the fraction of total bookings ultimately paid in each currency. Even if EUR comprises only a small percentage of daily transactions, that fraction can swing markedly from day to day, amplified by large booking volumes and exchange-rate shifts. Such changes directly affect net USD receipts and thus any revenue forecast that fails to anticipate compositional shifts may lead to significant errors. Moreover, once daily compositions are known, treasury teams can plan FX conversions more efficiently, hedge against unexpected fluctuations, or analyze how macroeconomic conditions influence guest payment preferences.
Each day’s currency proportions must lie in a simplex (they are nonnegative and sum to one), a constraint that complicates standard time-series methods. One classical approach is to use log-ratio transformations, as pioneered by Aitchison (1982) and extended in works such as Cargnoni et al. (1997), Ravishanker et al. (2001), and Mills (2010), among others. Transformed additive log-ratio VARMA (tVARMA) models have been fruitfully applied to compositional data (Brunsdon and Smith, 1998; Mills, 2009, 2010).
In contrast, the Dirichlet distribution models compositional data directly on the simplex. Benjamin et al. (2003) proposed frequentist Generalized ARMA models that could be extended to the simplex. Contributions to Dirichlet time series include Bayesian compositional state space models by Grunwald et al. (1993), da Silva et al. (2011), and da Silva and Rodrigues (2015); the Bayesian Dirichlet Auto-Regressive Moving Average (B-DARMA) model by Katz et al. (2024, 2025); and frequentist Dirichlet ARMA models by Zheng and Chen (2017).
Seminal works in Bayesian time series analysis include West (1996), Prado and West (2010), Barber et al. (2011), Berliner (1996), Pole et al. (2018), and Koop and Korobilis (2010). Bayesian methods have been applied to Vector Auto-Regression (VAR) and Vector Auto-Regression Moving Average (VARMA) models (Spencer, 1993; Uhlig, 1997; Bańbura et al., 2010; Karlsson, 2013). Bayesian generalized linear models (GLMs) for time series data, suitable for counts and categorical responses, have been explored by Brandt and Sandler (2012), Roberts and Penny (2002), Nariswari and Pudjihastuti (2019), Chen and Lee (2016), McCabe and Martin (2005), Berry and West (2020), Fukumoto et al. (2019), Silveira de Andrade et al. (2015), and West (2013).
The generalized auto-regressive conditional heteroskedasticity (GARCH) model (Bollerslev, 1986) is fundamental in modeling financial time series volatility. Extensions include non-normal likelihoods to capture heavy tails and skewness (Nelson, 1991; Engle, 2001). Bauwens et al. (2006) surveyed advancements in multivariate GARCH models.
Motivated by the need to capture both shifting means and evolving volatility in compositional time series—especially under sudden disruptions such as COVID-19—we develop a new class of models called Bayesian Dirichlet Auto-Regressive Moving Average with Dirichlet Auto-Regressive Conditional Heteroskedasticity (B-DARMA-DARCH). Unlike existing Dirichlet ARMA methods that assume a fixed or purely deterministic precision parameter, our framework introduces a Dirichlet Auto-Regressive Conditional Heteroskedasticity (DARCH) component. This addition endows the standard Bayesian Dirichlet ARMA model with a GARCH-like recursion on the precision $\phi_{t}$, enabling it to:
  1. 1.
    Model volatility clustering by letting $\phi_{t}$ depend stochastically on recent innovations and prior volatility, thereby reflecting abrupt regime changes and prolonged uncertainty.
  2. 2.
    Accommodate severe shocks (e.g., the onset of COVID-19) that induce persistent or recurrent surges in variance—phenomena not well captured by constant-precision or deterministic-variance models.
  3. 3.
    Preserve compositional constraints via the Dirichlet likelihood, ensuring forecasted currency shares remain valid proportions over time.
  4. 4.
    Leverage a full Bayesian framework to deliver predictive distributions (and credible intervals) that adapt automatically to time-varying volatility, resulting in more robust risk assessments and interval coverage.
These innovations allow B-DARMA-DARCH to simultaneously handle both systematic and stochastic changes in compositional data, providing a flexible, unified approach to dynamic heteroskedasticity on the simplex—a capability hitherto missing in Bayesian compositional time-series models.
We apply the B-DARMA-DARCH model to Airbnb’s currency fee proportions across different regions, demonstrating its effectiveness in capturing complex temporal dynamics and volatility patterns in real-world compositional data. We conduct simulation studies comparing the B-DARMA-DARCH model with the standard B-DARMA and Bayesian transformed VARMA (B-tVARMA) models. The B-DARMA-DARCH model consistently outperforms the other models in forecast accuracy and residual diagnostics, highlighting its robustness and applicability in scenarios with heteroskedasticity, structural breaks, and regime changes. By directly modeling compositional data within the simplex and incorporating time-varying volatility, our proposed B-DARMA-DARCH model offers a powerful tool for analyzing dynamic compositional time series.
The next section presents the B-DARMA-DARCH model. Section 3 presents simulation studies comparing the B-DARMA-DARCH to both the B-DARMA data model and a Bayesian transformed data normal VARMA model. Section 4 presents analysis of the Airbnb data. The paper closes with a short discussion.
2 A Bayesian Dirichlet Auto-Regressive Moving Average - Dirichlet Auto-Regressive Conditional Heteroskedasticity Model
2.1 DARMA Data model
Consider a $J$-component compositional time series $\{\mathbf{y}_{t}\}_{t=1}^{T}$ of length $T$ indexed by $t=1,\dots,T$, where each observation $\mathbf{y}_{t}=(y_{t1},\ldots,y_{tJ})^{\top}$ lies in the simplex
$\mathcal{S}^{J-1}=\left\{\mathbf{y}\in(0,1)^{J}:\sum_{j=1}^{J}y_{j}=1\right\}.$
We model $\mathbf{y}_{t}$ using the Dirichlet distribution with mean vector $\boldsymbol{\mu}_{t}$ and precision parameter $\phi_{t}>0$
$\mathbf{y}_{t}\mid\boldsymbol{\mu}_{t},\phi_{t}\sim\operatorname{Dirichlet}\left(\phi_{t}\boldsymbol{\mu}_{t}\right),$ (1)
where $\boldsymbol{\mu}_{t}=(\mu_{t1},\ldots,\mu_{tJ})^{\top}\in\mathcal{S}^{J-1}$ with probability density function (pdf)
$p(\mathbf{y}_{t}\mid\boldsymbol{\mu}_{t},\phi_{t})=\frac{\Gamma\left(\phi_{t}\right)}{\prod_{j=1}^{J}\Gamma\left(\phi_{t}\mu_{tj}\right)}\prod_{j=1}^{J}y_{tj}^{\phi_{t}\mu_{tj}-1},$ (2)
for $\mathbf{y}_{t}\in\mathcal{S}^{J-1}$.
2.1.1 Modeling the Mean using the Additive Log-Ratio Transformation
We model $\boldsymbol{\mu}_{t}$ as a function of prior observations $\mathbf{y}_{1},\ldots,\mathbf{y}_{t-1}$, prior means $\boldsymbol{\mu}_{1},\ldots,\boldsymbol{\mu}_{t-1}$ and known covariates $\boldsymbol{x}_{t}$ in a generalized linear model framework. As $\boldsymbol{\mu}_{t}$ and $\mathbf{y}$ are constrained, we model $\boldsymbol{\mu}_{t}$ after reducing dimension using the additive log ratio (alr) link
$\boldsymbol{\eta}_{t}\equiv{\textnormal{alr}}(\boldsymbol{\mu}_{t})=\left(\log\left(\frac{\mu_{t1}}{\mu_{tj^{*}}}\right),\ldots,\log\left(\frac{\mu_{tJ}}{\mu_{tj^{*}}}\right)\right)$ (3)
where $\boldsymbol{\eta}_{t}$ is the linear predictor, a $J-1$-vector taking values in $\mathbb{R}^{J-1}$, $j^{*}$ is a chosen reference component $1\leq j^{*}\leq J$, and the element of $\boldsymbol{\eta}_{t}$ that would correspond to the $j^{*}$th element $\log(\mu_{j^{*}}/\mu_{j^{*}})=0$ is omitted. Given $\boldsymbol{\eta}_{t}$, $\boldsymbol{\mu}_{t}$ is defined by the inverse of equation (3) where $\mu_{tj}=\exp(\eta_{tj})/(1+\sum_{j=1}^{J-1}\exp(\eta_{tj}))$ for $j=1,\dots,J$, $j\neq j^{*}$ and for $j=j*$, $\mu_{tj*}=1/(1+\sum_{j=1}^{J-1}\exp(\eta_{tj}))$.
The alr-transformed mean vector $\mathbf{\eta}_{t}$ is modeled as a Vector Auto-Regressive Moving Average (VARMA) process
$\mathbf{\eta}_{t}=\sum_{p=1}^{P}\boldsymbol{A}_{p}(\text{alr}(\mathbf{y}_{t-p})-\mathbf{X}_{t-p}\boldsymbol{\beta})+\sum_{q=1}^{Q}\boldsymbol{B}_{q}(\text{alr}(\mathbf{y}_{t-q})-\mathbf{\eta}_{t-q})+\mathbf{X}_{t}\boldsymbol{\beta},$ (4)
where $P$ is the VAR lag, $Q$ is the VMA lag, and $t=m+1,\dots,T$ where $m=\max(P,Q)$, $\boldsymbol{A}_{p}$ and $\boldsymbol{B}_{q}$ are $(J-1)\times(J-1)$, $\mathbf{X}_{t}\in\mathbb{R}^{J-1\times r_{\beta}}$ is a matrix of deterministic covariates, and $\boldsymbol{\beta}\in\mathbb{R}^{r_{\beta}}$ is a vector of regression coefficients.
2.1.2 Precision Parameter $\phi_{t}$
In the B-DARMA model, precision parameter $\phi_{t}$ is modeled with log link as a function of an $r_{\gamma}$-vector of covariates $\boldsymbol{z}_{t}$,
$\phi_{t}=\exp(\boldsymbol{z}_{t}\boldsymbol{\gamma}),$ (5)
where $\boldsymbol{\gamma}$ is an $r_{\gamma}$-vector of coefficients. If there are no covariates, $\log\phi_{t}=\gamma$ for all $t$.
2.2 DARCH Process for Precision Parameter $\phi_{t}$
We introduce a DARCH process for the precision parameter $\phi_{t}$, modeled with a log link
$\log(\phi_{t})$ $=\sum_{l=1}^{L}\alpha_{l}(\log(\phi_{t-l})-\mathbf{z}_{t-l}^{\prime}\boldsymbol{\gamma})+\sum_{k=1}^{K}\tau_{k}({\textnormal{alr}}(\mathbf{y}_{t-k})-\boldsymbol{\eta}_{t-k})^{\prime}({\textnormal{alr}}(\mathbf{y}_{t-k})-\boldsymbol{\eta}_{t-k})$
$\quad+\mathbf{z}_{t}^{\prime}\boldsymbol{\gamma},$ (6)
where $\alpha_{l}\in\mathbb{R}$ are auto-regressive coefficients for the log-precision, $\tau_{k}\in\mathbb{R}$ are coefficients associated with past innovations, and $\mathbf{z}_{t}\in\mathbb{R}^{r_{\gamma}}$ are known covariates with coefficients $\boldsymbol{\gamma}\in\mathbb{R}^{r_{\gamma}}$.
For ease of notation, we set $P=L$ and $Q=K$, but this assumption is easily relaxed as shown in section 3. The DARCH component models the volatility in the precision parameter $\phi_{t}$, addressing non-constant variability over time. By capturing volatility clustering and external influences, the DARCH process enhances the model’s ability to understand compositional data dynamics, improving forecasting and risk management.
Alternative log-ratio transformations, such as the centered log-ratio (CLR) or the isometric log-ratio (ILR), can also be used. A detailed discussion of these transformations is provided in the Supplementary Material (Section Alternative Log-Ratio Transformations).
2.3 Joint Predictive Distribution
Define the consecutive observations $\mathbf{y}_{a:b}=(\mathbf{y}_{a},\dots,\mathbf{y}_{b})^{\prime}$ for positive integers $a<b$. To be well defined, linear predictor (4) requires having $m$ previous observations $\mathbf{y}_{(t-m):(t-1)}$, and corresponding linear predictors $\boldsymbol{\eta}_{t-m},\ldots,\boldsymbol{\eta}_{t-1}$. In computing posteriors, we condition on the first $m$ observations $\mathbf{y}_{1:m}$ which then do not contribute to the likelihood. For the corresponding first $m$ linear predictors, on the right hand side of (4), we set $\boldsymbol{\eta}_{1},\ldots,\boldsymbol{\eta}_{m}$ equal to ${\textnormal{alr}}(\mathbf{y}_{1}),\ldots,{\textnormal{alr}}(\mathbf{y}_{m})$ which effectively omits the VMA terms $\boldsymbol{B}_{l}({\textnormal{alr}}(\mathbf{y}_{t-l}))$ from (4) and the MA terms $\tau_{l}$ $({\textnormal{alr}}(\mathbf{y}_{t-l})-\boldsymbol{\eta}_{t-l})^{\prime}({\textnormal{alr}}(\mathbf{y}_{t-l})-\boldsymbol{\eta}_{t-k})$ from (6) when $t-l\leq m$. In contrast, in (4) the VAR terms and $\boldsymbol{X}_{t}\boldsymbol{\beta}$ and in (6) the AR terms and $\mathbf{z}_{t}^{\prime}\boldsymbol{\gamma}$ are well defined for $t=1,\ldots,m$.
Define the $C$-vector $\boldsymbol{\theta}$ of all unknown parameters $\boldsymbol{\theta}=(\boldsymbol{A}_{prs},\boldsymbol{B}_{qrs},\boldsymbol{\beta}^{\prime},\boldsymbol{\gamma}^{\prime},\tau_{j},\alpha_{i})^{\prime}$, where $r,s=1,\ldots,J-1$ index all elements of matrices $\boldsymbol{A}_{p}$ and $\boldsymbol{B}_{q}$, $p=1,\ldots,P$, $q=1,\ldots,Q$, $i=1,\dots,L$, $j=1,\dots,K$ and $C=(P+Q)*(J-1)^{2}+L+K+r_{\beta}+r_{\gamma}$. Our prior beliefs about $\boldsymbol{\theta}$ ($p(\boldsymbol{\theta})$) are updated by Bayes’ theorem to give the posterior
$p(\boldsymbol{\theta}|\mathbf{y}_{1:T})=\frac{p(\boldsymbol{\theta})p(\mathbf{y}_{(m+1):T}|\boldsymbol{\theta},\mathbf{y}_{1:m})}{p(\mathbf{y}_{(m+1):T}|\mathbf{y}_{1:m})},$
[…]
2.4 Differences between B-DARMA, B-tVARMA, and B-DARCH
The differences among the Bayesian-ALR-transformed VARMA (B-tVARMA), B-DARMA, and B-DARCH models lie in their handling of compositional constraints and volatility. The B-tVARMA model transforms compositional data to an unconstrained space, which may introduce distortions if model assumptions like normality are violated. In contrast, the B-DARMA and B-DARCH models directly model data within the simplex using the Dirichlet distribution. For volatility, the B-tVARMA model assumes constant variance unless extended to include heteroskedasticity. The B-DARMA model allows a time-varying precision parameter $\phi_{t}$ modeled by exogenous factors but lacks a stochastic structure, potentially missing stochastic volatility patterns. The B-DARCH model introduces a dynamic $\phi_{t}$ with an ARMA structure, capturing both systematic and stochastic changes in volatility, such as clustering and sudden shifts. While the B-tVARMA model is simpler, it may inadequately represent compositional data and volatility dynamics. The B-DARMA model offers a balance but may miss important volatility features. The B-DARCH model, though more complex, comprehensively accounts for mean dynamics and volatility patterns, enhancing model fit and forecasting accuracy when heteroskedasticity is present.
3 Simulation Study
3.1 Overview
We conducted six simulation studies to evaluate the performance of the B-DARMA(1,0)-DARCH(1,1) model compared to the B-DARMA(1,0) and B-tVARMA(1,0) models. Each study generates data from either a DARMA, DARMA-DARCH, or tVARMA and introduces random shocks to simulate misreported observations (1-3) or redraws values for all parameters (4-6) to simulate a regime shift. Each study simulates 50 datasets, and evaluates out-of-sample forecast accuracy and in-sample residual autocorrelation.
3.2 Simulation Studies Setup
In all six studies, datasets with five compositional components were simulated over $T=100$ time points. The first 60 observations were used for training, and the remaining 40 for testing. For each simulation, a new set of parameters was drawn, giving different intercept vectors, auto-regressive matrices, and precision parameters across simulations.
All six simulation studies have the same underlying mean vector formulation for the latent ALR-transformed process
$\boldsymbol{\eta}_{t}=\boldsymbol{\beta}+\boldsymbol{A}(\text{alr}(\mathbf{y}_{t-1})-\boldsymbol{\beta}).$
The auto-regressive coefficient matrix $\boldsymbol{A}$ was a $4\times 4$ matrix with elements drawn from a uniform distribution, $A_{ij}\sim\mathcal{U}(-0.75,0.75)$, independently for each simulation. For each of the 50 simulations, the elements of the intercept vector $\boldsymbol{\beta}$ were generated by drawing five elements $\beta_{j}^{*},j=1,\dots,5$ from $\mathcal{N}(0.2,0.03^{2})$, then normalized to sum to 1 ($\beta_{j}^{*}=\frac{\beta_{j}^{*}}{\sum_{j^{*}=1}^{5}\beta_{j}^{*}})$, transformed using the ALR transformation to give the $4\times 1$ intercept vector $\boldsymbol{\beta}=\text{alr}(\boldsymbol{\beta}^{*})$. For each simulation, the initial values were
$y_{1j}=.2+\epsilon_{j},\ \epsilon_{j}\sim N(0,.01^{2}),\ j=1,\dots,5$
followed by $y_{1}=\frac{y_{1}}{\sum y_{1}}.$
3.2.1 Simulation Studies 1 & 4: DARMA DGP
In Simulations 1 and 4, the data-generating process (DGP) was a DARMA model, $\mathbf{y}_{t}|\mathbf{\mu}_{t},\phi_{t}\sim\text{Dirichlet}(\phi_{t}\mathbf{\mu}_{t})$ with a constant precision parameter $\log(\phi_{t})=\phi_{0}$. The precision parameter $\phi_{0}$ was drawn independently for each simulation from a uniform distribution between 6 and 7.5, $\phi_{0}\sim\mathcal{U}(6,7.5).$
3.2.2 Simulation Studies 2 & 5: DARCH DGP
The data in the second and fifth simulations were generated using a DARMA-DARCH process, where $\phi_{t}$ was generated
$\log(\phi_{t})=\phi_{0}+\alpha(\log(\phi_{t-1})-\phi_{0})+\tau\sum_{i=1}^{4}(\text{alr}(y_{t-1,i})-\eta_{t-1,i})^{2}.$
The true intercept $\phi_{0}$ was drawn independently for each simulation from a uniform distribution between 6 and 7.5, $\phi_{0}\sim\mathcal{U}(6,7.5).$ We fixed parameters $\alpha=0.8$ and $\tau=-0.95$ to simulate data with high volatility persistence and strong responsiveness to past innovations, creating heteroskedasticity patterns that test the B-DARCH model’s ability to capture dynamic volatility.
3.2.3 Simulation Studies 3 & 6: tVARMA DGP
The data in the third and sixth simulations were generated using a tVARMA process with
$\boldsymbol{\eta}_{t}\sim\mathcal{N}(\boldsymbol{\beta}+\boldsymbol{A}(\text{alr}(\mathbf{y}_{t-1})-\boldsymbol{\beta}),\sigma^{2}\boldsymbol{\Sigma})$
$\mathbf{y}_{t}=\text{alr}^{-1}(\boldsymbol{\eta}_{t}).$
A single precision parameter $\sigma$ was drawn independently for each simulation from $\sigma\sim\mathcal{U}(0.05,0.5).$ The covariance matrix $\boldsymbol{\Sigma}$ was generated to be positive-definite with complex interdependencies $\boldsymbol{\Sigma}=\mathbf{M}^{\top}\mathbf{M},$ where $\mathbf{M}$ was a $4\times 4$ matrix with elements drawn independently from $\mathcal{U}(-0.3,0.3)$ for each simulation.
3.2.4 Simulated Shocks
In simulations 1-3, random shocks were introduced into the training data to simulate misreported observations. Time points were selected by sampling intervals from a Poisson(6) distribution and cumulatively summing them. At these times $l$, $\mathbf{y}_{t}$ was replaced with a random composition generated by independently drawing each component from a Uniform(0,1) distribution and normalizing the vector to sum to 1,
$y_{lj}\sim\mathcal{U}(0,1),\ \mathbf{y}_{\text{misreported}[l]}=\frac{y_{lj}}{\sum_{j=1}^{5}y_{lj}},j=1,\dots,5.$
3.2.5 Regime Shifts
In Simulation 4-6, we introduced regime shifts at a random time point $t_{\text{shift}}$ uniformly selected from [10, 50] by redrawing the intercept vector $\boldsymbol{\beta}$, the auto-regressive matrix $A$ and the precision parameters (either $\phi_{0}$, $\sigma$, $\Sigma$) for the DARMA/DARCH models or tVARMA model. For $t=t_{\text{shift}}$ to $t=t_{\text{shift}}+9$, data were generated using these new parameters, simulating a regime shift in the data-generating process. After this period, parameters reverted to their original values, allowing us to evaluate the models’ robustness to temporary structural changes.
3.3 Model Comparisons
We fitted three models to each simulated dataset for comparison: the B-DARMA(1,0)-DARCH(1,1), the Bayesian transformed VARMA (B-tVARMA(1,0)), and the B-DARMA(1,0) models. All models shared the following priors: the intercepts $\boldsymbol{\beta}$ were given $\text{Normal}(0,0.3^{2})$ priors, and the elements of the auto-regressive matrices $\boldsymbol{A}_{p}$ had $\text{Normal}(0,1)$ priors.
For the B-DARMA(1,0)-DARCH(1,1) model, the initial precision parameter $\phi_{0}$ was assigned a $\text{Normal}(7,1.5^{2})$ prior, and the AR and MA terms for $\phi_{t}$ had $\text{Normal}(0.35,0.5^{2})$ and $\text{Normal}(-0.75,0.5^{2})$ priors, respectively.
[…]
All models were fit using Stan with four chains of 1,000 iterations each (500 warm-up), yielding 2,000 posterior samples.
3.4 Metrics
We assess model performance using Forecast Root Mean Squared Error (FRMSE) and Forecast Mean Absolute Error (FMAE) for each of the five components. These metrics quantify the out-of-sample forecast accuracy
${\text{FRMSE}}_{j}=\left(\frac{1}{40}\sum_{t=61}^{100}\left(y_{tj}-\bar{\mu}_{tj}\right)^{2}\right)^{\frac{1}{2}}$
${\text{FMAE}}_{j}=\frac{1}{40}\sum_{t=61}^{100}\left|y_{tj}-\bar{\mu}_{tj}\right|$
where $\bar{\mu}_{tj}$ represents the posterior mean of the predicted value at time $t$ for component $j$. We then summarize by averaging them across all five components ${\text{FRMSE}}_{mean}=\frac{1}{5}\sum_{j=1}^{5}{\text{FRMSE}}_{j}$ and ${\text{FMAE}}_{mean}=\frac{1}{5}\sum_{j=1}^{5}{\text{FMAE}}_{j}$.
Additionally, we assess the Partial Autocorrelation Function (PACF) (see Supplementary Section Definition of the Partial Autocorrelation Function (PACF)) of the sums of squared standardized residuals to evaluate how well each model captures residual dynamics during the in-sample period. For each model, residuals are standardized according to the model’s underlying distribution (see Supplementary Section Standardized Residuals).
3.5 Results
3.5.1 Simulation Studies 1-3: Shocks
The left side of table 1 shows the forecast performance FRMSE and FMAE metrics for the B-DARMA, B-DARCH, and B-tVARMA models across 50 simulated datasets in three different scenarios.
Across simulation studies 1–3, the B-DARCH model consistently had the lowest FRMSE and FMAE values compared to the second-best model, the B-tVARMA. In simulation study 1, FRMSE and FMAE were 1.6% and 1.4% lower, respectively. In simulation study 2, they were 12.6% and 7.7% lower. In simulation study 3, FRMSE and FMAE were .8% and 3.0% lower than B-tVARMA.
Figure S1 shows the mean PACF values across all simulations for B-DARCH, B-DARMA, and B-tVARMA models. The mean PACF values for the B-DARCH model consistently remain closer to zero across the majority of lags, with fewer pronounced PACF values. Notably, the B-DARMA and B-tVARMA models show large negative PACF values, particularly for early lags (lags 1-5). These large values indicate that residuals in these models have substantial autocorrelation, suggesting less efficient capture of the underlying dependencies across time within the data. In contrast, B-DARCH shows minimal autocorrelation for early lags, reflecting an overall improvement in residual behavior and a better fit to the true data-generating process.
3.5.2 Simulation Studies 4-6: Regime Shifts
The right side of Table 1 shows the forecast performance metrics FRMSE and FMAE for the B-DARMA, B-DARCH, and B-tVARMA models across 50 simulations with regime shifts. The B-DARCH model consistently had the lowest FRMSE and FMAE values, outperforming the second-best model (B-tVARMA) with 14.6% lower FRMSE and 8.4% lower FMAE in simulation study 4, 4.7% and 4.6% lower in simulation study 5, and 10.5% and 4.8% lower in simulation study 6. These results highlight the B-DARCH model’s comparative advantage in adapting to sudden changes in the underlying dynamics.
Figure S2 shows the mean PACF values for simulations 4-6 for B-DARCH, B-DARMA, and B-tVARMA models. The B-DARCH model has positive peaks for early lags, with autocorrelation generally settling closer to zero across subsequent lags. Notably, the magnitude of the PACF values for B-DARCH is smaller than those of the other models, except for Simulation 4, where the B-DARMA model shows a smaller magnitude at lag 1.
In contrast, both B-DARMA and B-tVARMA have strong positive PACF values for the first few lags, which then gradually decrease to negative values at later lags, indicating greater persistence of autocorrelation effects. This pattern suggests that B-DARMA and B-tVARMA models struggle more with handling shifts in the underlying process, resulting in larger autocorrelations in early lags compared to the B-DARCH model.
These results indicate that the B-DARCH model’s dynamic volatility component enables it to adapt more effectively to structural breaks, capturing both systematic and stochastic changes in volatility. In contrast, the B-DARMA and B-tVARMA models, lacking a stochastic volatility structure, were less effective in handling the regime shifts, leading to higher forecast errors.
Model Simulation 1 (DARMA Shock) Simulation 4 (DARMA Regime Shift)
Average FRMSE Average FMAE Average FRMSE Average FMAE
B-DARMA 38.6 2.94 25.3 2.58
B-DARCH 38.0 2.87 18.1 2.19
B-tVARMA 38.6 2.91 21.2 2.37
Model Simulation 2 (DARCH Shock) Simulation 5 (DARCH Regime Shift)
Average FRMSE Average FMAE Average FRMSE Average FMAE
B-DARMA 11.2 1.84 20.4 2.28
B-DARCH 9.79 1.69 18.3 2.09
B-tVARMA 11.3 1.83 19.2 2.19
Model Simulation 3 (tVARMA Shock) Simulation 6 (tVARMA Regime Shift)
Average FRMSE Average FMAE Average FRMSE Average FMAE
B-DARMA 25.6 2.43 32.5 2.59
B-DARCH 24.1 2.24 22.9 2.18
B-tVARMA 24.3 2.31 25.7 2.29
Table 1: Summary of Model Performance Metrics times 100 on test set ($T=40$) across 50 simulations for Simulation Studies 1–6. Average FRMSE is the average Forecasted Root Mean Squared Error across all five components; Average Forecasted Mean Absolute Error is the average FMAE across all five components.
4 Airbnb Data Analysis
The dataset comprises daily records from January 1, 2017, to December 31, 2020, detailing the proportion of fees in U.S. Dollars (USD) originating from various billing currencies. The data is categorized into four anonymized diverse geographic regions, referred to as Regions 1-4. Within each region, the fees are broken down by the top five currencies specific to that region, with the remaining currencies grouped into an “other” category. The top five currencies for each region are fixed for the entire period. The training period spans from January 1, 2017, to June 30, 2020. The subsequent three months (July 1, 2020 to September 30, 2020) are used for validation and model specification, while the final 3 months (October 1, 2020, to December 31, 2020) constitute the test set.
4.1 Exploratory Data Analysis
4.1.1 Regional Currency Dynamics
The currency proportions for each region are shown in Figures 1 and Supplementary Figures S3 and S4. Using the U.S. Dollar (USD) as the reference category, 30-day rolling ALR means are shown in Supplementary Figures S5 and S6 while 30-day rolling ALR variances are shown in Supplementary Figure S7 and Figure 2. The correlations of the ALR values with their lag 1 values is shown in Supplementary Figure S8.
Across all four regions, currencies exhibit seasonal cycles and trends relative to USD. Regions 1 and 3 display pronounced yearly seasonality for multiple currencies, with Region 4 exhibiting the strongest seasonal patterns. Pre-COVID-19 (before 2020), currencies in Regions 1 and 2 generally showed steady levels or consistent declines relative to USD, while in Region 3, the Australian Dollar (AUD) and other currencies displayed distinct upward trends. During the COVID-19 period, starting around January or February of 2020, significant shifts are seen: in Region 1, all currencies except the Canadian Dollar (CAD) saw large declines in their means at the expense of the USD; in Region 2, the Euro (EUR) and Swiss Franc (CHF) experienced the largest increases as the USD decline; in Region 3, AUD continued its sharp rise, while the New Zealand Dollar (NZD) initially declined but rapidly recovered; in Region 4, the Brazilian Real (BRL) experienced sharp increases in its rolling mean before being overtaken by USD, and the Mexican Peso (MXN) peaked before a drastic decline.
Lag 1 correlations show strong autocorrelations within currencies, with high self-correlations observed in EUR and GBP in Region 1 and in BRL, Chilean Peso (CLP), and MXN in Region 4. Positive lag 1 cross correlations were found among currencies in Region 2, except for CAD, while in Region 3, NZD and AUD were strongly correlated, and EUR was negatively correlated with other currencies. In Region 4, negative correlations between BRL and EUR, as well as between CLP and EUR, suggest opposing movements between these currency pairs.
Figure 2 and Supplementary Figure S7 show the 30-day rolling variance of each currency’s ALR-transformed proportions. Rather than a single abrupt change, we observe recurrent episodes of elevated variability, often persisting for several weeks before subsiding. For instance, Region 1 (top left) shows multiple surges for EUR and GBP, culminating in early 2020 and tapering mid-year, while Region 2 (top right) begins with a high-variance phase (notably in CAD and CHF), settles to moderate levels around 0.1–0.2, and then rises again at the onset of 2020. Region 3 (bottom left) experiences shorter bursts around 2018 and late 2019, each followed by another peak in early 2020, suggesting that once volatility increases, it may linger rather than immediately revert. Finally, Region 4 (bottom right) transitions from moderate day-to-day swings to abrupt climbs in BRL and CLP near 2019, with additional spikes in BRL and MXN emerging after 2020. In every region, these heightened-variance intervals emerge at distinct times and exhibit varying durations, indicating that volatility in currency shares is shaped by more than one-off shocks, and can arise repeatedly under evolving market conditions. Such patterns underscore the importance of accommodating dynamic, recurrent shifts in variance for effective modeling and forecasting.
Figure 1: Airbnb data analysis - proportion of fees by billing currency for four regions from Jan 1, 2017 to June 30, 2020. AUD is the Australian dollar, BRL is the Brazillian Real, CAD is the Canadian Dollar,CHF is the Swiss Franc, CLP is the Chilean Peso,EUR is the European Euro, GBP is the Great British Pound, MXN is the Mexican Peso, NZD is the New Zealand Dollar, and USD is the US Dollar.
Figure 2: Airbnb data analysis - proportion of fees by currency- 30-day rolling ALR variance for four regions from Jan 1, 2017 to Dec 31, 2019. USD is the reference component.
4.2 Objective
The primary goal of this analysis is to compare the performance of three model classes: Bayesian Dirichlet Auto-Regressive Moving Average (B-DARMA), Bayesian transformed Vector Auto-Regressive Moving Average (B-tVARMA), and Bayesian Dirichlet Auto-Regressive Moving Average with Dirichlet Auto-Regressive Conditional Heteroskedasticity (B-DARMA-DARCH).
4.3 Comparison Models
The B-DARCH and B-DARMA models share the same model for $\boldsymbol{\eta}$ but differ in their treatment of the precision parameter. In the B-DARCH model, the precision parameter has its own AR and MA components, allowing for dynamic heteroskedasticity that evolves over time. In contrast, the B-DARMA model does not include these components in the precision parameter, relying solely on auto-regressive and moving average terms in the mean structure to account for time dependence.
The third model, B-tVARMA, is a Bayesian normal VARMA model applied to transformed data. It transforms the observed multivariate time series $\mathbf{y}_{t}$ using the additive log-ratio (alr) transformation, resulting in ${\textnormal{alr}}(\mathbf{y}_{t})\in\mathbb{R}^{J-1}$. The transformed data is then modeled as ${\textnormal{alr}}(\mathbf{y}_{t})\sim N(\boldsymbol{\eta}_{t},\mathbf{\Sigma})$, where $\boldsymbol{\eta}_{t}$ has time-varying mean defined in (4), and $\bf{\Sigma}$ is an unknown $J-1\times J-1$ positive definite covariance matrix. The covariance matrix $\bf{\Sigma}$ is modeled through its Cholesky factor $L_{\Sigma}$, scaled by $\sigma$, $\bf{\Sigma}$ $=\sigma^{2}\Omega$.
All models are fit with STAN (Stan Development Team, 2022) in R. We run $4$ chains with $2000$ iterations each with a warm up of $1000$ iterations for a posterior sample of $4000$. Initial values for all parameters are selected randomly from the interval $[-1,1]$.
4.4 Seasonal and Trend Terms
Figure 1 and Supplementary figure S3 show varying levels of yearly and weekly seasonality in the currency data in different regions. There is a recurring yearly pattern driven by popular travel periods and a clear weekly seasonality reflecting variations between days of the week. Consequently, we include weekly and yearly seasonal variables and a trend variable in the predictors $\mathbf{X}_{t}$ and the same weekly and yearly seasonal variables in $\boldsymbol{z}_{t}$. Both $\phi_{t}$ and the elements of $\boldsymbol{\eta}_{t}$ are modeled with a linear trend and Fourier terms for seasonal variables, using pairs of $\left(\sin{\frac{2k\pi t}{w_{\text{season}}}},\cos{\frac{2k\pi t}{w_{\text{season}}}}\right)$ for $k=1,\dots,K_{\text{season}}\leq\frac{w_{\text{season}}}{2}$, with $w_{\text{season}}=7$ for weekly seasonality and $w_{\text{season}}=365.25$ for yearly seasonality. The orthogonality of the Fourier terms aids in model convergence. The B-DARCH, B-DARMA, and B-tVARMA all share the same $\mathbf{X}_{t}$ while the B-DARCH and B-DARMA share the same $\boldsymbol{z}_{t}$.
4.5 Model Specification and Validation
The models are fit using data from January 1, 2017, to June 30, 2020, as the training set. The validation set, spanning July 1, 2020, to September 30, 2020, is used to determine optimal lags and exogenous covariates. The test set from October 1, 2020, to December 31, 2020, is used for final model comparison.
For the initial specification, we fix $P=1$, $Q=0$, and $K_{\text{week}}=3$ for each region, allowing the intra-year Fourier terms to vary. The optimal number of Fourier terms ($K_{\text{year}}$) is selected based on the model’s performance on the validation set. With the optimal ($K_{\text{year}}$) established, we vary $P$ and $Q$ to identify the best-performing orders of the auto-regressive (AR) and moving average (MA) components on the validation set. Using these specifications, the models are retrained up until September 30, 2020, and tested on the independent test set for the B-DARMA, B-DARMA-DARCH, and B-tVARMA models.
4.6 Priors for B-DARCH, B-DARMA, and B-tVARMA
For all models (B-DARCH, B-DARMA, and B-tVARMA), we assume each component’s auto-regressive (AR) and moving average (MA) elements, $a_{rs}$ and $b_{rs}$, are strongest with their own lags. Thus, we assign weakly informative normal priors $N(.4,.5^{2})$ to $a_{rrp},b_{rrq}$ for diagonal elements ($r=1,\dots,4,q=1,\dots,3$), and $N(0,.5^{2})$ for non-diagonal elements in the $\boldsymbol{A}$ and $\boldsymbol{B}$ matrices.
The priors for the Fourier, trend, and intercept regression coefficients are $N(0,1)$, $N(0,.1^{2})$, and $N(0,2^{2})$ respectively. For the B-DARMA and B-DARCH, the precision model predictors $\boldsymbol{z}_{t}$ have the same seasonal terms as the mean models, with corresponding priors for the coefficients in $\boldsymbol{\gamma}$ matching those of the $\boldsymbol{\eta}$ terms. The B-DARCH has a $N(0,1)$ prior on its AR $\alpha$ and MA $\tau$ terms.
For the B-tVARMA, we impose a half normal prior $N^{+}(0,.5^{2})$ on the scale parameter $\sigma$ and the correlation matrix $\Omega$ has a $\text{LKJ}(3)$ prior on its Cholesky factor $L_{\Omega}$.
4.7 Results
4.7.1 Model Specification
Out-of-sample validation results (Supplementary Table S1) show that Region 1 performed best with $K_{\text{year}}=8$ and $(P,Q)=(1,0)$, yielding a average FMAE of 0.0036 and total FRSS of .0173. Region 2 achieved optimal performance with $K_{\text{year}}=10$ and $(P,Q)=(1,1)$, resulting in a average FMAE of .0127 and tRSS of .2018. In Region 3, the optimal model had $K_{\text{year}}=10$ and $(P,Q)=(1,1)$, with a average FMAE of .0283 and total FRSS of 1.1395. Region 4 was best modeled with $K_{\text{year}}=8$ and $(P,Q)=(3,2)$, with a FMAE of .0116 and FRSS of .1553.
4.7.2 Performance on Test Set
The Forecast Mean Absolute Error (FMAE) and Forecast Residual Sum of Squares (FRSS) for each region and major currency are shown in Table 2, while Figures 36 display the corresponding out‐of‐sample forecasts $\bar{\mu}_{tj}$. When errors are aggregated over all currencies within each region, the B-DARCH model consistently delivers the lowest total FMAE and FRSS. For example, in Region 1 the total FMAE (FRSS) for B-DARCH is 2.84 (3.15), which is markedly lower than 7.65 (19.11) for B-DARMA and 5.20 (10.62) for B-tVARMA. In Region 2, the advantage of B-DARCH is even more pronounced: its total FMAE is 13.55 (with an average of 2.26) and total FRSS is 73.93 (average 12.32), compared to 14.71 (2.45) and 81.87 (13.65) for B-DARMA, and 17.13 (2.86) and 104.43 (17.41) for B-tVARMA. Regions 3 and 4 exhibit similar aggregate advantages for B-DARCH.
At the individual-currency level, however, there are still several cases in which either B-DARMA or B-tVARMA yield lower FMAE—particularly in Regions 3 and 4. In Region 2, by contrast, the B-DARCH yields the smallest FMAE across all major currencies (CAD, CHF, EUR, GBP, USD, and “other”). In Regions 3 and 4, there remain a few currencies that favor one of the alternative models. For instance, in Region 3, B-DARMA produces slightly lower FMAE than B-DARCH for JPY (2.17 vs. 2.20) and NZD (2.91 vs. 3.57), while in Region 4, B-DARMA outperforms B-DARCH for EUR (0.58 vs. 1.29) but not for most of the other currencies. Notably, B-tVARMA is especially competitive for certain currencies in Region 1 (e.g., EUR), likely because its assumption of constant variance aligns well with periods of relatively stable volatility. Nonetheless, when the full basket of currencies is considered, the time-varying volatility framework of B-DARCH leads to the most accurate overall forecasts.
Supplementary Table S2 shows the empirical coverage rates of the 95% credible intervals for the forecasts across regions and currencies. The B-DARCH model exhibits relatively stable coverage that is generally closer to the nominal 95% level, with mean coverage rates of 0.92, 0.91, 0.87, and 0.91 in Regions 1, 2, 3, and 4, respectively. By contrast, the B-DARMA model shows greater variability, with mean coverages ranging from 0.82 in Region 1 down to 0.55 in Region 3, and 0.82 in Region 2 and 0.84 in Region 4 B-tVARMA tends to undercover, with mean rates of 0.63, 0.86, 0.69, and 0.85 across Regions 1–4. These results suggest that although B-tVARMA can occasionally yield competitive point forecasts, its assumption of time-invariant volatility often leads to overconfident interval estimates. In contrast, the dynamic volatility specification in B-DARCH not only improves point forecast accuracy but also delivers more reliable interval estimates.
Region Currency FMAE FRSS
B-DARCH B-DARMA B-tVARMA B-DARCH B-DARMA B-tVARMA
1 CAD 0.55 0.87 1.21 0.49 1.07 1.75
1 EUR 0.61 2.19 0.55 0.53 5.23 0.55
1 GBP 0.18 0.43 0.21 0.05 0.22 0.06
1 JPY 0.09 0.18 0.06 0.01 0.04 0.01
1 USD 1.08 3.13 2.52 1.91 11.70 7.65
1 other 0.33 0.85 0.66 0.17 0.85 0.60
1 total 2.84 7.65 5.20 3.15 19.11 10.62
1 average 0.47 1.27 0.87 0.53 3.19 1.77
2 CAD 0.35 0.38 0.37 0.22 0.24 0.27
2 CHF 0.83 0.89 0.96 0.97 1.14 1.36
2 EUR 4.51 4.85 5.51 29.70 34.50 42.20
2 GBP 4.06 4.13 5.32 23.00 24.00 35.30
2 USD 1.28 1.68 1.68 3.04 4.59 4.60
2 other 2.52 2.78 3.29 17.00 17.40 20.70
2 total 13.55 14.71 17.13 73.93 81.87 104.43
2 average 2.26 2.45 2.86 12.32 13.65 17.41
3 AUD 6.29 8.24 7.78 57.50 88.50 80.80
3 EUR 0.63 0.93 0.87 0.52 1.09 0.99
3 JPY 2.20 2.17 2.34 23.70 24.10 22.70
3 NZD 3.57 2.91 3.57 14.60 9.96 14.70
3 USD 0.82 0.79 1.02 1.00 0.98 1.52
3 other 2.69 5.09 3.06 13.30 34.80 15.30
3 total 16.20 20.12 18.64 110.62 159.43 136.01
3 average 2.70 3.35 3.11 18.44 26.57 22.67
4 BRL 2.84 4.36 6.38 10.80 23.60 46.20
4 CLP 2.32 2.35 3.09 7.76 8.45 13.00
4 EUR 1.29 0.58 0.64 2.08 0.51 0.60
4 MXN 1.27 2.85 1.54 2.36 8.71 2.98
4 USD 2.70 4.14 2.94 8.68 19.60 9.93
4 other 0.79 0.97 1.94 0.87 1.46 4.93
4 total 11.21 15.25 16.53 32.55 62.33 77.64
4 average 1.87 2.54 2.76 5.43 10.39 12.94
Table 2: Airbnb data analysis: Forecast Mean Absolute Error (FMAE) and Forecast Residual Sum of Squares (FRSS) for Regions 1 to 4 and top currencies. FMAE and FRSS values are multiplied by $10^{2}$.
Figure 3: Airbnb data analysis - Region 1: 92-day forecasts (blue) with 95% confidence intervals, represented by the shaded regions around the forecast lines, for each of the six currencies, from October 1, 2020, to December 31, 2020, compared to the actual values (red) observed over the preceding six months. The models compared include B-DARCH, B-DARMA, and B-tVARMA.
Figure 4: Airbnb data analysis - Region 2: 92-day forecasts (blue) with 95% confidence intervals, represented by the shaded regions around the forecast lines, for each of the six currencies, from October 1, 2020, to December 31, 2020, compared to the actual values (red) observed over the preceding six months. The models compared include B-DARCH, B-DARMA, and B-tVARMA.
Figure 5: Airbnb data analysis - Region 3: 92-day forecasts (blue) with 95% confidence intervals, represented by the shaded regions around the forecast lines, for each of the six currencies, from October 1, 2020, to December 31, 2020, compared to the actual values (red) observed over the preceding six months. The models compared include B-DARCH, B-DARMA, and B-tVARMA.
Figure 6: Airbnb data analysis - Region 4: 92-day forecasts (blue) with 95% confidence intervals, represented by the shaded regions around the forecast lines, for each of the six currencies, from October 1, 2020, to December 31, 2020, compared to the actual values (red) observed over the preceding six months. The models compared include B-DARCH, B-DARMA, and B-tVARMA.
4.7.3 Residual Analysis
The partial autocorrelation function (PACF) of the sum of standardized squared residuals across all components on the test set was computed for each region and model (Supplementary Figures S9S12). All models exhibit a large peak at lag 1, reflecting immediate persistence in the residuals. However, the magnitude of this spike is noticeably smaller for the B-DARCH model. We consider a PACF value to be significant if it exceeds a threshold of 0.2; any autocorrelation values above this level indicate meaningful residual dependence.
Beyond lag 1, the B-DARCH model shows only one PACF value exceeding the 0.2 significance threshold, suggesting effective capture of short-term dependencies and minimal residual autocorrelation at higher lags. In contrast, both B-DARMA and B-tVARMA have PACF values surpassing the 0.2 threshold at multiple lags beyond lag 1, particularly between lags 2 and 10, indicating that residual dependencies persist across a broader range of lags in these models.
These suggest that the B-DARCH model achieves better residual independence and stability, with remaining residual structure potentially explained by unmodeled market-specific holidays, events, and policy implementations.
4.7.4 Parameter Estimates.
A complete discussion of the posterior densities for the DARCH parameters (auto-regressive $\alpha$ and moving average $\tau$ coefficients) in each region appears in the Supplementary Material (Section Parameter Estimates). Briefly, Regions 2 and 3 exhibit higher $\alpha$ values indicating more persistent volatility, while Region 1 shows moderate persistence ($0.750.25<\alpha<$ 0.75), and Region 4 displays a more complex multi-lag structure. These region-specific patterns underscore the flexibility of our B-DARMA-DARCH approach to accommodate various volatility regimes.
5 Discussion
We introduced the Bayesian Dirichlet Auto-Regressive Moving Average with Dirichlet Auto-Regressive Conditional Heteroskedasticity (B-DARMA-DARCH) model, a novel approach for modeling and forecasting dynamic compositional time series with time-varying volatility. By incorporating a DARCH component into the B-DARMA framework, the model effectively captures both mean dynamics and heteroskedasticity inherent in compositional data.
Simulation studies demonstrated the superior performance of the B-DARMA-DARCH model across six scenarios, including data generated from DARMA, DARCH, and tVARMA processes with random shocks simulating misreported observations and regime changes. The B-DARMA-DARCH model consistently achieved the lowest forecast error metrics and exhibited minimal residual autocorrelation, indicating its robustness in capturing complex volatility dynamics, especially in the presence of heteroskedasticity.
Applying the B-DARMA-DARCH model to Airbnb’s currency fee proportions across different regions, we showed that it effectively captures temporal dynamics and volatility patterns in real-world data. The model generally outperformed the standard B-DARMA and B-tVARMA models in terms of forecast accuracy and residual diagnostics. Notably, the inclusion of the DARCH component allowed the model to account for significant disruptions, such as those caused by the COVID-19 pandemic, highlighting its robustness in the face of structural breaks and extreme events.
While the B-DARMA-DARCH model introduces additional complexity due to the dynamic modeling of the precision parameter, this complexity is justified by the improved model fit, empirical coverage rates, and forecasting accuracy in the presence of heteroskedasticity. The model offers a comprehensive framework that accounts for both mean dynamics and volatility patterns, which is particularly important in financial applications where volatility significantly influences decision-making and risk assessment.
However, the computational demands are higher, necessitating careful tuning of sampling parameters to ensure convergence and efficiency. Future research could focus on optimizing computational strategies or exploring approximate inference methods to mitigate these demands.
Several limitations warrant consideration. The current model models each region independently, even though some currencies are shared across multiple regions. This assumption of independence may not hold in practice, as economic events affecting a shared currency could simultaneously impact multiple regions. Incorporating hierarchical structures or accounting for inter-regional dependencies could enhance model performance by capturing these cross-regional correlations, albeit at the cost of increased computational complexity. Additionally, extending the model to handle zero-inflated compositional data or integrating external covariates that influence both the mean and volatility components could broaden its applicability.
Acknowledgement
The authors thank Sean Wilson, Jess Needleman, Jenny Cheng, Erica Savage, Liz Medina, and Yuling Kuo for helpful discussions, Ellie Mertz, Adam Liss, and Peter Coles for championing the research, and Lauren Mackevich for their indispensable operational support.
Supplementary Materials
Alternative Log-Ratio Transformations
Although we have used the additive log-ratio (ALR) link in equations (3), (4), and (5), other transformations to and from the simplex may be employed. A common choice is the centered log-ratio (clr), defined as
${\textnormal{clr}}(\mathbf{y})\;=\;\Bigl{(}\,\ln\bigl{(}\tfrac{y_{1}}{g(\mathbf{y})}\bigr{)},\;\ln\bigl{(}\tfrac{y_{2}}{g(\mathbf{y})}\bigr{)},\;\dots,\;\ln\bigl{(}\tfrac{y_{J}}{g(\mathbf{y})}\bigr{)}\Bigr{)},$
where $g(\mathbf{y})=(\,y_{1}\,y_{2}\,\cdots\,y_{J})^{1/J}$ is the geometric mean of $\mathbf{y}$. The clr mapping places all $J$ components in a transformed $\mathbb{R}^{J}$ space but introduces the constraint that their sum is zero.
Another option is to use the Isometric Log-Ratio (ilr) transformation. For each index $j$, the ilr-transformed coordinate can be expressed as
$z_{j}=\sqrt{\frac{r_{j}}{r_{j}+1}}\,\ln\biggl{(}\frac{g_{j}(\mathbf{y)}}{\bigl{(}\prod_{i\in H_{j}}y_{i}\bigr{)}^{1/r_{j}}}\biggr{)},$
where $g_{j}(\mathbf{y})$ is defined as the geometric mean of a chosen subset $S_{j}\subseteq\mathbf{y}$, and $H_{j}$ is the complement of $S_{j}$ in $\mathbf{y}$. The value $r_{j}$ is the number of elements in $H_{j}$. The particular way in which $S_{j}$ and $H_{j}$ are selected depends on the specific application and how one wishes to interpret the ilr coordinates.
These mappings are mutually related by linear transformations (Egozcue et al., 2003) and each provides a bijection from the $(J-1)$-dimensional simplex onto a $(J-1)$-dimensional subset of $\mathbb{R}^{J}$ or $\mathbb{R}^{J-1}$. Therefore, if the same priors and model assumptions are consistently adapted to each log-ratio space, the resulting compositional models remain equivalent—differing only in how one embeds the data in a Euclidean domain.
Definition of the Partial Autocorrelation Function (PACF)
The PACF at lag $k$ measures the partial correlation between the SSR at time $t$ and the SSR at time $t-k$, controlling for the effects of intermediate lags $1$ to $k-1$. It is defined as the last coefficient $\phi_{kk}$ in the autoregressive model of order $k$
$\text{SSR}_{t}=\phi_{k1}\text{SSR}_{t-1}+\phi_{k2}\text{SSR}_{t-2}+\dots+\phi_{k,k-1}\text{SSR}_{t-(k-1)}+\phi_{kk}\text{SSR}_{t-k}+\varepsilon_{t}$ (7)
where $\varepsilon_{t}$ is the white noise error term at time $t$. The coefficient $\phi_{kk}$ represents the PACF at lag $k$ and can be calculated recursively using the Yule-Walker equations
$\phi_{kk}=\frac{r_{k}-\sum_{j=1}^{k-1}\phi_{k-1,j}r_{k-j}}{1-\sum_{j=1}^{k-1}\phi_{k-1,j}r_{j}}$
Here, $r_{k}$ is the autocorrelation at lag $k$, and $\phi_{k-1,j}$ are the PACF coefficients from the previous lag $k-1$.
By examining the PACF of the SSRs, we can identify the presence of residual autocorrelation at various lags, which indicates whether the model has adequately captured the temporal dependencies in the data.
Standardized Residuals
For the B-DARMA and B-DARCH models, which are based on the Dirichlet distribution as defined in Section 2, we compute standardized residuals using the standard deviations derived from the Dirichlet parameters.
At each time t, the residuals are calculated as $\text{residual}_{tj}=y_{tj}-\mu_{tj},\quad\text{for }j=1,\dots,J.$ The variance of each component $y_{tj}$ under the Dirichlet distribution, Dirichlet($\phi_{t}\boldsymbol{\mu}_{t}$), is $\operatorname{Var}[y_{tj}]=\frac{\mu_{tj}(1-\mu_{tj})}{\phi_{t}+1}$, with standard deviation $\sigma_{tj}=\sqrt{\frac{\mu_{tj}(1-\mu_{tj})}{\phi_{t}+1}}.$
[…]
By examining the PACF of $\{\text{SSR}\}_{t=1}^{T}$ we assess the presence of autocorrelation in the residuals. A well-specified model should produce residuals with minimal autocorrelation, indicating that temporal dependencies have been adequately captured.
Parameter Estimates
Supplementary Figure S13 shows the posterior densities of the auto-regressive ($\alpha$) and moving average ($\tau$) coefficients for each region. In Region 1, the posterior of $\alpha$ is concentrated between 0.25 and 0.75, peaking around 0.45, indicating moderate volatility persistence; shocks have a noticeable but not prolonged effect on volatility. In Region 2, the $\alpha$ coefficient has posterior mean around 0.8, reflecting strong dependence of current volatility on past volatility, and the $\tau$ coefficient centers around $-2.75$, indicating that large deviations from expected currency proportions significantly increase / decrease future volatility—characteristic of volatility clustering. Region 3 shows $\alpha$ tightly concentrated between 0.85 and 0.90 and $\tau$ centered around $-1.2$ suggesting very strong volatility persistence. In contrast, Region 4 exhibits complex dynamics: $\alpha_{1}$ is centered around $0.15$, indicating small persistence from the first lag; $\alpha_{2}$ is skewed positive and centered around $0.35$, suggesting a stronger influence from the second lag; $\alpha_{3}$ peaks slightly above 0.5, indicating moderate persistence; and the $\tau$ parameter is concentrated near $0.02$, implying minimal influence of past innovations.
Overall, the higher $\alpha$ values in Regions 2 and 3 indicate that volatility in currency fee proportions is more persistent in these regions. The negative $\tau$ coefficients suggest that large deviations—such as those caused by abrupt travel restrictions, policy changes, and fluctuating foreign exchange (FX) rates lead to increased future volatility. This implies that Regions 2 and 3 are more susceptible to external shocks affecting travel and currency markets, resulting in prolonged volatility in currency fee proportions. In contrast, Region 1, with moderate volatility persistence, tends to return to stability more rapidly after shocks, possibly due a more stable currency environment or a larger share of domestic travel. Region 4’s dynamics, characterized by less emphasis on immediate past volatility and innovations, may be attributed to regional factors such as diversified travel sources, differing policy responses, or economic structures.
Figure S1: Average PACF of the sum of squared standardized residual values across all 50 simulations for studies 1-3. B-DARMA-DARCH consistently exhibits lower PACF values beyond lag 1, indicating reduced residual autocorrelation and suggesting that its dynamic precision component better captures volatility shocks in these simulated compositional time series.
Figure S2: Average PACF of the sum of squared standardized residual values across all 50 simulations for studies 4-6. The B-DARMA-DARCH model (red) again shows smaller autocorrelation at higher lags, whereas B-DARMA (green) and B-tVARMA (red) leave more persistent structures in the residuals. This result underlines the advantage of modeling time-varying volatility when faced with abrupt changes in compositional data.
Figure S3: Airbnb data analysis- Proportion of fees by currency for four regions: weekly seasonal behavior. AUD is the Australian dollar, BRL is the Brazillian Real, CAD is the Canadian Dollar,CHF is the Swiss Franc, CLP is the Chilean Peso,EUR is the European Euro, GBP is the Great British Pound, MXN is the Mexican Peso, NZD is the New Zealand Dollar, and USD is the US Dollar. Strong weekly cycles appear in many currencies, motivating the inclusion of weekly seasonal terms in our model’s mean structure.
Figure S4: Airbnb data analysis - proportion of fees by billing currency for four regions from Jan 1, 2020 to June 30, 2020. AUD is the Australian dollar, BRL is the Brazillian Real, CAD is the Canadian Dollar,CHF is the Swiss Franc, CLP is the Chilean Peso,EUR is the European Euro, GBP is the Great British Pound, MXN is the Mexican Peso, NZD is the New Zealand Dollar, and USD is the US Dollar. Distinct spikes and dips across multiple currencies (e.g., sharp increase in BRL or EUR) reflect significant disruptions.
Figure S5: Airbnb data analysis - proportion of fees by currency- 30-day rolling ALR means for four regions from Jan 1, 2020 to June 30, 2020. USD is the reference component.
Figure S6: Airbnb data analysis - proportion of fees by currency- 30-day rolling ALR means for four regions from Jan 1, 2017 to June 30, 2020. USD is the reference component. Apart from seasonal patterns, we see notable multi-year trends and cyclical behaviors (e.g., increasing share of EUR in Region 2, upward drift in AUD in Region 3), underlining the presence of both trend and seasonality in compositional dynamics.
Figure S7: Airbnb data analysis - proportion of fees by currency- 30-day rolling ALR variance for four regions from Jan 1, 2017 to June 30, 2020. USD is the reference component.The data display periods of high variance that sometimes persist for weeks, demonstrating “volatility clustering.” These episodes justify the GARCH-like specification in our B-DARMA-DARCH model, which allows the precision parameter $\phi_{t}$ to track spikes in uncertainty over time.
Figure S8: Airbnb data analysis- correlation between ALR and lagged ALR values for each currency in the four regions. USD is the reference component. High positive autocorrelations in many currencies (e.g., EUR, GBP) highlight the strong temporal persistence. Negative cross-lag correlations for some pairs (e.g., EUR vs. BRL in Region 4) indicate that certain currencies move in opposite directions. Both effects reinforce the need for a multi-component time-series framework with auto- and cross-dependence.
Figure S9: Airbnb data analysis: Partial Autocorrelation Function (PACF) values for the sum of squared standardized residuals (SSR) on the test set from October 1, 2020, to December 31, 2020, for Region 1.
Figure S10: Airbnb data analysis: Partial Autocorrelation Function (PACF) values for the sum of squared standardized residuals (SSR) on the test set from October 1, 2020, to December 31, 2020, for Region 2.
Figure S11: Airbnb data analysis: Partial Autocorrelation Function (PACF) values for the sum of squared standardized residuals (SSR) on the test set from October 1, 2020, to December 31, 2020, for Region 3.
Figure S12: Airbnb data analysis: Partial Autocorrelation Function (PACF) values for the sum of squared standardized residuals (SSR) on the test set from October 1, 2020, to December 31, 2020, for Region 4.
Figure S13: Airbnb data analysis: Densities of DARCH model coefficients for all 4 regions. Region 1 was modeled with a DARCH(1,0), Regions 2-3 modeled with a DARCH(1,1), and Region 4 modeled with a DARCH(3,1). Each plot shows the estimated density of a specific $\alpha_{l}$ (ar_phi) or $\tau_{k}$ (ma_phi) parameter. Regions 2 and 3 exhibit high $\alpha$ indicating persistent volatility once spikes occur, whereas Region 4’s coefficients suggest moderate or delayed response
Supplementary Figures and Tables
P Q Fourier FMAE FRSS
1 0 6 0.38 1.75
1 0 8 0.36 1.73
1 0 10 0.45 2.65
1 0 12 6.40 442.25
1 0 14 18.66 3662.84
1 0 16 12.42 1638.80
1 0 18 18.87 3809.00
0 1 8 5.81 404.83
1 1 8 1.05 14.60
1 2 8 0.98 15.02
1 3 8 1.03 16.12
2 1 8 0.40 3.96
2 2 8 6.57 716.43
2 3 8 2.09 60.38
3 1 8 8.17 984.45
3 2 8 8.91 1209.30
3 3 8 1.17 16.30
(a) Region 1
P Q Fourier FMAE FRSS
1 0 6 2.15 49.95
1 0 8 2.10 48.46
1 0 10 1.76 38.08
1 0 12 6.46 343.58
1 0 14 13.31 1621.76
1 0 16 9.23 832.27
1 0 18 13.97 1892.17
0 1 10 4.74 193.91
1 1 10 1.27 20.18
1 2 10 1.28 23.08
1 3 10 2.81 48.12
2 1 10 3.08 263.17
2 2 10 2.43 91.08
2 3 10 2.38 73.18
3 1 10 5.84 344.45
3 2 10 2.67 106.77
3 3 10 5.88 380.70
(b) Region 2
P Q Fourier FMAE FRSS
1 0 6 3.55 149.23
1 0 8 3.49 144.21
1 0 10 3.40 136.33
1 0 12 5.78 360.13
1 0 14 11.49 1281.57
1 0 16 9.37 862.47
1 0 18 12.80 1541.94
0 1 10 10.79 982.42
1 1 10 2.83 113.95
1 2 10 3.22 153.84
1 3 10 3.19 150.57
2 1 10 2.89 226.67
2 2 10 3.12 148.21
2 3 10 2.98 131.34
3 1 10 3.94 222.45
3 2 10 3.55 190.52
3 3 10 3.42 166.84
(c) Region 3
P Q Fourier FMAE FRSS
1 0 6 4.72 216.17
1 0 8 4.64 212.57
1 0 10 4.92 236.31
1 0 12 8.00 490.97
1 0 14 12.05 1058.80
1 0 16 9.57 750.15
1 0 18 12.15 1148.46
0 1 8 9.60 693.31
1 1 8 2.66 76.77
1 2 8 1.63 31.03
1 3 8 1.86 38.07
2 1 8 1.43 47.14
2 2 8 1.50 26.37
2 3 8 1.49 24.93
3 1 8 1.33 20.92
3 2 8 1.16 15.53
3 3 8 1.90 40.07
(d) Region 4
Table S1: Airbnb data analysis: B-DARCH model performance across 4 regions – average Forecast Mean Absolute Error (FMAE) and total Forecast Residual Sum of Squares (FRSS) for the top 6 currencies for the validation set from July 1, 2020 to Sep 30, 2020. FMAE and FRSS values are multiplied by $10^{2}$. The table identifies the optimal $(P,Q)$ and number of Fourier terms for each region, showing how model complexity affects forecasting accuracy. Lower FMAE/FRSS indicate better predictive fit, guiding final model selection for each region.
Table S2: Coverage rates (%) of 95% credible intervals for each region and currency under B-DARCH, B-DARMA, and B-tVARMA. The mean coverage is computed across currencies in each region. The B-DARMA-DARCH coverage tends to be closest to the nominal rate across regions (e.g., 92% in Region 1, 95% in Region 4), whereas B-tVARMA often undercovers due to fixed variance assumptions, and B-DARMA coverage can be inconsistent when volatility spikes.
Region Currency B-DARCH B-DARMA B-tVARMA
1 CAD 0.85 0.88 0.72
EUR 0.83 0.60 0.65
GBP 0.96 0.69 0.56
JPY 1.00 0.94 0.82
USD 0.90 0.94 0.44
other 0.96 0.84 0.57
Mean 0.92 0.82 0.63
2 CAD 0.913 0.913 0.783
CHF 0.967 0.891 0.924
EUR 0.957 0.913 0.967
GBP 0.870 0.772 0.924
USD 0.859 0.641 0.663
other 0.880 0.761 0.870
Mean 0.91 0.82 0.86
3 AUD 0.81 0.29 0.43
EUR 0.93 0.77 0.79
JPY 0.71 0.73 0.78
NZD 0.83 0.33 0.43
USD 0.95 0.81 0.81
other 0.97 0.39 0.89
Mean 0.87 0.55 0.69
4 BRL 0.97 0.87 0.97
CLP 0.69 0.54 0.30
EUR 0.97 1.00 1.00
MXN 0.98 0.90 0.99
USD 0.89 0.74 0.96
other 1.00 0.99 0.90
Mean 0.91 0.84 0.85